OpenStudy (anonymous):

Given f(x)=2x+3 and g(x)=x2is exponent) + 5, find (f o g) (x) ?

OpenStudy (anonymous):

f(g(x))=2(x^2+5)+3 f(g(x))=2x^2+10+3 f(g(X))=2x^2+13

OpenStudy (anonymous):

u got it

OpenStudy (anonymous):

medal plz

OpenStudy (anonymous):

f(x) = x^2 - 6x+4 using the discriminant